nean 64 and od omens heights ds between 70 % to 16. 50 people are selected randomly from a certain population and it is found that 12 people in the sample are over 6 feet tall. a) what is the point estimate of the proportion of people in the population who are over 6 feet tall? (2pts) b) Construct a 99% confidence interval for the mean over 6 feet tall. (6pts) elected ey had at leas ga95% opulation who have t least pts) c) Interpret the 99% confidence interval for the mean. (2pts)
Answer #1

a) Point estimate = 12/50 = 0.24

b) E=2.58*\sqrt{\frac{0.24*0.76}{50}}

= 0.16

Lower bound = 0.24-0.16 = 0.08

Upper Bound = 0.24+0.16 = 0.40

c) we are 99% confident that proportion of people who are over 6 feet tall will lie in this interval